OpenStudy (allymariex3):

Solve. x2 – 3x = 18 A.x = –3 and x = 6 B.x = 3 and x = –6 C.x = 2 and x = –3 D.x = –2 and x = 3

OpenStudy (anonymous):

x^2-3x-18=0 x^2 - 6x +3x -18=0 x(x-6)+3(x-6)=0 thus we have (x-3)(x-6)=0 thus x=6,3

OpenStudy (anonymous):

isnt it X2=3X+18 like that?